OpenStudy (anonymous):

How do I write an equation of a line that goes through point (4, –5) and is parallel to line y = 2x + 8?

OpenStudy (anonymous):

Parallel lines have equal slope

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Now, y=2x+c , where c is the constant. will be the equation of the required line.

OpenStudy (anonymous):

(4, –5) use this point to find the value of c

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

Not quite.

OpenStudy (anonymous):

y=2x+c Now, Since the line passes through point (4,-5) -5=2*4 +c c=-13

OpenStudy (anonymous):

Now, substitute c in y=2x+c and get the equation of the line

OpenStudy (anonymous):

Ooooh, I see. Thank you!

OpenStudy (anonymous):

:)yw